OpenStudy (anonymous):

What is the solution set of |2x – 3| – 8 = –1?

OpenStudy (cwrw238):

|2x – 3| – 8 = –1 |2x - 3| = 7 the expression in the | | can be positive or negative so 2x - 3 = 7 giving 2x = 10 and x = 5 also 2x - 3 = -7 giving x = -2 solution set is (-2,5)

OpenStudy (anonymous):

i got (-1.5.3)

OpenStudy (cwrw238):

what are your 2 elements? -1 and 5.3 ?